Nine college students had eaten the following number of times at a fast food restaurant for dinner in the last ten days:

8, 6, 5, 4, 7, 0, 10, 10, 5

Find the mean, median, mode, and range for these data. Please show all of your work.

Answer by CubeyThePenguin(3113)

ordered: 0, 4, 5, 5, 6, 7, 8, 10, 10

mean = (8 + 6 + 5 + 4 + 7 + 0 + 10 + 10 + 5)/9 = 6.11

median = middle number = 6

mode = 5 and 10 (both occur twice)

range = (highest) - (lowest) = 10 - 0 = 10